OpenStudy (zubhanwc3):

use substitution to evaluate the indefinite integral. solve for x in this substitution.

OpenStudy (zubhanwc3):

\[\int\limits_{}^{}((x+1)\sqrt{2-x})dx\]

OpenStudy (zubhanwc3):

i found the indefinite integral to be \[2/5(2-x)^{3/2}(x+3) + C\]

OpenStudy (zubhanwc3):

i dont know what to do, to solve for x

OpenStudy (anonymous):

umm.. from what i know, if you have found the integral, you don't need to solve for x nor can not

OpenStudy (anonymous):

You're missing a negative sign but in general you would break it up to: \[\int x \sqrt{2-x}dx + \int \sqrt{2-x}dx\] The second one is easy, the first one can be solved with integration by parts.

OpenStudy (zubhanwc3):

would my answer be that, or do i have to solve it T_T

OpenStudy (zubhanwc3):

without intervals, i dont know wat to do T_T

OpenStudy (amoodarya):